Find the eigen value and eigen vector of A = [ 3 3 1 5 ] 


1
Expert's answer
2021-08-19T17:25:27-0400
"A=\begin{bmatrix}
 3 & 3 \\
 1 & 5
\end{bmatrix}"

"\det(A-\lambda I)=\begin{vmatrix}
 3-\lambda & 3 \\
 1 & 5-\lambda
\end{vmatrix}=0"

"( 3-\lambda)( 5-\lambda)-3(1)=0"

"15-8\lambda+\lambda^2-3=0"

"\lambda^2-8\lambda+12=0"

"(\lambda-2)(\lambda-6)=0"

The eigenvalues are "\lambda_1=2, \lambda_2=6."


"\lambda=2"


"A-\lambda I=\begin{bmatrix}
 3-\lambda & 3 \\
 1 & 5-\lambda
\end{bmatrix}=\begin{bmatrix}
 1 & 3 \\
 1 & 3
\end{bmatrix}"

"R_2=R_2-R_1"


"\begin{bmatrix}
 1 & 3 \\
 0 & 0
\end{bmatrix}"

Solve


"\begin{bmatrix}
 1 & 3 \\
 0 & 0
\end{bmatrix}\begin{bmatrix}
 x_1 \\
 x_2
\end{bmatrix}=\begin{bmatrix}
 0 \\
 0
\end{bmatrix}"

If we take "x_2=t," then "x_1=-3t."

The eigenvector is "\vec x=\begin{bmatrix}
 -3 \\
 1
\end{bmatrix}"



"\lambda=6"


"A-\lambda I=\begin{bmatrix}
 3-\lambda & 3 \\
 1 & 5-\lambda
\end{bmatrix}=\begin{bmatrix}
 -3 & 3 \\
 1 & -1
\end{bmatrix}"

"R_1=-R_1\/3"


"\begin{bmatrix}
 1 & -1 \\
 1 & -1
\end{bmatrix}"


"R_2=R_2-R_1"


"\begin{bmatrix}
 1 & -1 \\
 0 & 0
\end{bmatrix}"

Solve


"\begin{bmatrix}
 1 & -1 \\
 0 & 0
\end{bmatrix}\begin{bmatrix}
 y_1 \\
 y_2
\end{bmatrix}=\begin{bmatrix}
 0 \\
 0
\end{bmatrix}"

If we take "y_2=t," then "y_1=t."

The eigenvector is "\vec y=\begin{bmatrix}
 1 \\
 1
\end{bmatrix}"


Eigenvalue: "2," eigenvector: "\begin{bmatrix}
 -3 \\
 1
\end{bmatrix}."

Eigenvalue: "6," eigenvector: "\begin{bmatrix}
 1 \\
 1
\end{bmatrix}."